OpenStudy (anonymous):

word problem: Kristin spent $131 on shirts.Fancy shirts cost $28 and plain shirts cost $15.If she bought a total of 7 then how many of each kind did she buy?

OpenStudy (anonymous):

f + p = 7 and 28f + 15p = 131. Solvable system. Re-write the 1st equation into f = 7 - p and substitute into the second equation. You will get your value for "p". Go back with this value to the first equation and you will get "f".